jacksono6:

In 1990, the price of a gallon of unleaded gasoline was $1.37. In 1998, it was $1.02 per gallon. What is the decrease expressed as a percent change? Round to the nearest tenth of a percent if necessary.

jimthompson5910:

Here's a similar problem The price of gas starts off at $3.00 one year, and then the next year it drops to $2.50 What is the percentage decrease?


Solution: A = old value = 3.00 B = new value = 2.50 C = percent change C = [(B-A)/A]*100% C = [(2.50-3)/3]*100% C = (-0.50/3)*100% C = (-0.1667)*100% C = -16.67% The negative percentage means we have a percent decrease. If C were positive, then it would be a percent increase. The solution to this example problem is roughly 16.67% decrease Use these steps to solve your question.